# Learning Effective Representations for Retrieval Using Self-Distillation with Adaptive Relevance Margins

### v1

@@ 5.5. Comparison to Baselines
To compare the effectiveness of our proposed approach to established alternatives, Table 4 lists nDCG@10 and Recall@1000 scores in the full ranking setup for each of our three loss variants, as well as scores for state-of-the-art bi-encoder models trained via teacher distillation by Hofstätter et al. (2021). We also include evaluation scores of the reproduction study by Wang and Zuccon (2023). Both studies finetune the distilBERT model exclusively, therefore we restrict our comparison to distilBERT-based runs. The table is grouped into approaches relying on self-distillation (adaptive and distributed targets), teacher distillation (MarginMSE, original (Hofstätter et al., 2021) and reproduced (Wang and Zuccon, 2023)), and no distillation (static targets and BM25).To compare our approach to teacher distillation, we obtained run files from the reproduction study by Wang and Zuccon (2023), and tested for statistical equivalence of our approaches against their best-performing run (paired TOST, , , [TREC-DL19], [TREC-DL20], Bonferroni correction applied per test group). Statistical equivalence is indicated in Table 4 using a symbol. No run files were available for the study by Hofstätter et al. (2021), hence no statistical testing could be carried out here. To compare data efficiency, the table also includes the number of training triplets seen by each model. The early-stopping criterion of Hofstätter et al. (2021) terminates runs after 700-800k steps at a batch size of 32. This puts a conservative estimate of the data seen during training for their models at least at 22.4M training triplets.Wang and Zuccon (2023) terminate earlier, at steps with the same batch size for their highest-scoring run, i.e., at 14.6M triplets seen. Their other runs use similar amounts of data. Training time is not comparable due to the different software and hardware setups used between our experiments and prior work. Yet, all teacher distillation approaches require an additional cross-encoder forward pass for the full dataset of triplets for batch sampling.For nDCG@10, the best self-distillation approach (distributed) scores 0.07 lower than the best original teacher-distillation approach (TAS-B batch sampling with in-batch negatives). To the scores attained in the reproduction study, a difference of at most 0.04 can be observed, and the effectiveness is statistically equivalent in all cases. For Recall@1000, a similar trend is observable: compared to the original teacher-distillation approach, ours scores lower by 0.06 at most (adaptive with in-batch negatives), yet still statistically equivalent to the reproduction study. This attests to our approach’s competitive effectiveness, matching the scores of teacher-distillation-based approaches. When comparing the static target approach, we can make two observations: first, tuning the target hyperparameter yields a large increase in Recall@1000, but less so for nDCG@10. Second, it also matches the effectiveness of the teacher-distilled models in the reproduction study, further calling into question if the expense of teacher score inference is worth the marginal increase in effectiveness. Yet, the static targets variant necessitates hyperparameter tuning, and thus incurs higher computation cost (see Table 3) while offering similar effectiveness as self-distillation. The data efficiency of the self-distillation approach is between 2x/3x (distributed) and 7x/11x (adaptive) better than the teacher-based approaches.
@@ 6. Conclusion
In this paper, we investigated self-distillation with distributed relevance margins as a new approach to fine-tuning bi-encoder models for retrieval and ranking tasks. We proposed adaptive relevance margins as a novel self-distillation loss function. In contrast to previous teacher-distillation-based approaches, it eliminates the need for a teacher model and batch sampling procedure, and is hyperparameter-free. This is achieved by replacing the teacher signal with a text-to-text similarity score estimated by the fine-tuned model itself. Additionally, we introduced distributed targets, extending the adaptive approach with a novel approach to using in-batch information. Both variants perform implicit in-batch negative mining.Our experimental study demonstrated the effectiveness and training efficiency of the self-distillation approach. We have shown that both adaptive and distributed targets achieve competitive effectiveness while requiring substantially less training triplets and training compute time compared to state-of-the-art teacher-distillation methods. At the same time, we have demonstrated that teacher-distillation only offers marginal yields in effectiveness over hyperparameter-tuned static margin losses. Our results underline the viability of self-distillation as an alternative for training bi-encoder models, as it is data-efficient and hyperparameter-free.Self-distillation could also find applications in information retrieval beyond bi-encoder training. In future work we will investigate if it can be applied to pairwise cross-encoder training, which also uses triplets of a query and a positive/negative document. Further, the implicit negative mining aspect of adaptive and distributed targets can be extended further, e.g., by incorporating different scaling-, weighting-, or cutoff-schemes. Optimizing a single positive-negative pair against a distribution of target margins is also a promising direction for future work that we will investigate.
@@ Acknowledgements.
This publication has been partially supported by the ScaDS.AI Center for Scalable Data Analytics and Artificial Intelligence, funded by the Federal Ministry of Education and Research of Germany and by the Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus, by a Research Fellowship for Harrisen Scells from the Alexander von Humboldt Foundation, and by the OpenWebSearch.eu project, funded by the European Union (GA 101070014). We would also like to thank Shuai Wang for providing us with run files and additional insight about of their reproduction efforts (Wang and Zuccon, 2023).References
@@
- Karpukhin et al. (2020) Vladimir Karpukhin, BarlasOguz, Sewon Min, Patrick S. H. Lewis,Ledell Wu, Sergey Edunov,Danqi Chen, and Wen-tau Yih.2020. Dense Passage Retrieval for Open-Domain QuestionAnswering. In Proceedings of the 2020 Conferenceon Empirical Methods in Natural Language Processing, EMNLP 2020, Online,November 16-20, 2020, B. Webber,T. Cohn, Y. He, andY. Liu (Eds.). ACL,6769–6781.
@@
- Sharma et al. (2024) Rohan Sharma, Kaiyi Ji,zhiqiang xu, and Changyou Chen.2024. AUC-CL: A Batchsize-Robust Framework forSelf-Supervised Contrastive Representation Learning. InThe Twelfth International Conference on LearningRepresentations.

### v2

@@ 5.5. Comparison to TAS-B
To compare the effectiveness of our proposed approach to established alternatives, Table 4 lists nDCG@10 and Recall@1000 scores in the full ranking setup for each of our three loss variants, as well as scores for state-of-the-art bi-encoder models trained via teacher distillation by Hofstätter et al. (2021). We also include evaluation scores of the reproduction study by Wang and Zuccon (2023). Both experiments fine-tune the distilBERT model exclusively, therefore we restrict our comparison to distilBERT-based runs. The table is grouped into approaches relying on self-distillation (adaptive and distributed targets), teacher distillation (MarginMSE, original (Hofstätter et al., 2021) and reproduced (Wang and Zuccon, 2023)), and no distillation (static targets and BM25). As before, for tuned static targets, TREC-DL 19 is deliberately used both for hyperparameter optimization and to report results.To compare our approach to teacher distillation, we obtained run files from the reproduction study by Wang and Zuccon (2023), and tested for statistical equivalence of our approaches against their best-performing run (paired TOST, , , [TREC-DL 19], [TREC-DL 20], Bonferroni correction applied per test group). Statistical equivalence is indicated in Table 4 using a symbol. No run files were available for the experiments by Hofstätter et al. (2021), hence no statistical testing could be carried out here. To compare data efficiency, the table also includes the number of training triplets seen by each model. The early-stopping criterion of Hofstätter et al. (2021) terminates runs after 700-800k steps at a batch size of 32. This puts a conservative estimate of the data seen during training for their models at least at 22.4M training triplets.Wang and Zuccon (2023) terminate earlier, at with the same batch size for their highest-scoring run, i.e., at 14.6M triplets seen. Their other runs use similar amounts of data. Training time is not comparable due to the different software and hardware setups used between our experiments and prior work. Yet, all teacher distillation approaches require an additional cross-encoder forward pass for the full dataset of for batch sampling.
| System | TCV | NFC | NQ | HQA | FQA | AgA | T20 | CDS | QUO | DBP | SCD | FEV | CFE | SCF | I.-D. | Z.-S. | ||
| Ours | Adaptive (w/o. IB) | |||||||||||||||||
| Static (Untuned, w. IB) | ||||||||||||||||||
| Adaptive (w. IB) | ||||||||||||||||||
| Static (Untuned, w/o IB) | ||||||||||||||||||
| Distributed (w/o E.S.) | ||||||||||||||||||
| Static (Tuned, w. IB) | ||||||||||||||||||
| Static (Tuned, w/o IB) | ||||||||||||||||||
| Distributed (w. E.S.) | ||||||||||||||||||
| Dns./Lex. | DPR | |||||||||||||||||
| ANCE | ||||||||||||||||||
| BM25 | ||||||||||||||||||
| TAS-B | ||||||||||||||||||
| SOTA | ColBERT | |||||||||||||||||
| Contriever | ||||||||||||||||||
| SPLADE | ||||||||||||||||||
For nDCG@10, the best self-distillation approach (distributed) scores 0.07 lower than the best original teacher distillation approach (TAS-B batch sampling with in-batch negatives). To the scores attained in the reproduction study, a difference of at most 0.04 can be observed, and the effectiveness is statistically equivalent in all cases. For Recall@1000, a similar trend is observable: compared to the original teacher distillation approach, ours scores lower by 0.06 at most (adaptive with in-batch negatives), yet still statistically equivalent to the reproduction study. This attests to our approach’s competitive effectiveness, matching the scores of teacher distillation-based approaches. When comparing the static target approach, we can make two observations: first, tuning the target hyperparameter yields a large increase in Recall@1000, but less so for nDCG@10. Second, it also matches the effectiveness of the teacher-distilled models in the reproduction study, further calling into question if the expense of teacher score inference is worth the marginal increase in effectiveness. Yet, the static targets variant necessitates hyperparameter tuning (see Table 3). The data efficiency of the self-distillation approach is between 2x/3x (distributed) and 7x/11x (adaptive) better than the teacher-based approaches.
5.6. Comparison on BEIR
We evaluate on the public portion of the BEIR benchmark (Thakur et al., 2021) and comparing to baseline and SOTA scores from the official leaderboard (Kamalloo et al., 2024). Our comparison systems include: BM25 as a lexical baseline; three dense approaches using in-batch negatives (DPR (Karpukhin et al., 2020), ANCE (Xiong et al., 2021), and TAS-B (Hofstätter et al., 2021)); and three SOTA systems (ColBERT (Khattab and Zaharia, 2020), Contriever (Izacard et al., 2022), and SPLADE (Formal et al., 2021)). Table 5 presents these results. For fair comparison with TAS-B, we use finetuned distilBERT variants, though MPNet variants showed increased effectiveness ().The results demonstrate that our proposed approaches achieve competitive performance in in-domain settings while requiring substantially lower computational and data resources than established methods. Specifically the distributed and tuned static margins variants exhibit superior effectiveness, exceeding the score of TAS-B, and nearly being on-par with current SOTA models. Adaptive margins perform slightly worse, but still comptetitive.Out-of-domain evaluation reveals patterns similar to in-domain results: distributed and tuned static margins demonstrate superior effectiveness, while adaptive margins perform slightly worse, particularly without in-batch negatives. Parameter tuning for static margins consistently improves effectiveness. However, out-of-domain effectiveness relative to other dense retrievers remains lower than in-domain, confirming the observation of Thakur et al. (2021) that datasets with strong domain shift are problematic for dense retrievers, particularly evident on Touche-2020, FiQA, and CQADupStack.The distributed approach performs competitively against the best-in-class teacher-based TAS-B method with an average difference of only 0.09 in the zero-shot setting, and exceeding it by 0.2 in the in-domain setting. Overall, it outperforms DPR and closely matching ANCE and TAS-B at substantially less data and training computation. Early stopping successfully diminishes overfitting, yielding the same in-domain at improved out-of-domain effectiveness. Using MPNet as backbone model, average scores increase by equaling ANCE, and nearly matching TAS-B (paired TOST, , , ). This positions distributed margins primarily as an highly effective method for retrieval finetuning in in-domain low-resource settings, at diminished but still competitive zero-shot capabilities.
@@ 6. Conclusion
We introduced self-distillation with adaptive and distributed relevance margins for bi-encoder fine-tuning in retrieval tasks. Our approach introduces two key innovations: a self-distillation mechanism that leverages the model’s own similarity assessments, and a distributed target scheme that exploits batch-level information. These eliminate three common requirements in retrieval model training: teacher models, specialized batch sampling, and hyperparameter tuning. Our empirical evaluation demonstrates that distributed targets achieve competitive effectiveness, being statistically equivalent on the same dataset to state-of-the-art teacher distillation methods. As it offers very strong in-domain capabilities, our approach is particularly valuable for environments with limited evaluation data or compute resources. Our results further reveal that traditional teacher distillation offers only marginal improvements over hyperparameter-tuned static margins, questioning the computational cost-benefit ratio of complex teacher architectures. This establishes self-distillation as a practical alternative, especially for purpose-built in-domain models and low-resource settings, since it offers on-par effectiveness with a substantially reduced amount of required data and compute cost. It could find applications in information retrieval beyond bi-encoder training, e.g., extending to pairwise cross-encoder training. While we focus on web text ranking as target task, our proposed approach only relies on a suitable embedding space and could thus be transferred to other modalities, e.g., product, music, or image search.
@@ Acknowledgements.
This work has been partially funded by the German Federal Ministry of Research, Technology, and Space (BMFTR) under Grant № 01IS24077B; by the ScaDS.AI Center for Scalable Data Analytics and Artificial Intelligence, funded by the BMFTR and by the Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus under Grant № ScaDS.AI; by a Research Fellowship for Harrisen Scells from the Alexander von Humboldt Foundation; and by the OpenWebSearch.eu project, funded by the European Union (GA 101070014). We thank Shuai Wang for providing us with run files and additional insight about of their reproduction efforts (Wang and Zuccon, 2023).References
@@
- Formal et al. (2021) Thibault Formal, Benjamin Piwowarski, and Stéphane Clinchant. 2021. SPLADE: Sparse Lexical and Expansion Model for First Stage Ranking. In SIGIR ’21: The 44th International ACM SIGIR Conference on Research and Development in Information Retrieval, Virtual Event, Canada, July 11-15, 2021, Fernando Diaz, Chirag Shah, Torsten Suel, Pablo Castells, Rosie Jones, and Tetsuya Sakai (Eds.). ACM, 2288–2292. doi:10.1145/3404835.3463098
@@
- Izacard et al. (2022) Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2022. Unsupervised Dense Information Retrieval with Contrastive Learning. Trans. Mach. Learn. Res. 2022 (2022).
@@
- Kamalloo et al. (2024) Ehsan Kamalloo, Nandan Thakur, Carlos Lassance, Xueguang Ma, Jheng-Hong Yang, and Jimmy Lin. 2024. Resources for Brewing BEIR: Reproducible Reference Models and Statistical Analyses. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2024, Washington DC, USA, July 14-18, 2024, Grace Hui Yang, Hongning Wang, Sam Han, Claudia Hauff, Guido Zuccon, and Yi Zhang (Eds.). ACM, 1431–1440. doi:10.1145/3626772.3657862
- Karpukhin et al. (2020) Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick S. H. Lewis, Ledell Wu, Sergey Edunov, Danqi Chen, and Wen-tau Yih. 2020. Dense Passage Retrieval for Open-Domain Question Answering. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing, EMNLP 2020, Online, November 16-20, 2020, Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, 6769–6781. doi:10.18653/V1/2020.EMNLP-MAIN.550
- Khattab and Zaharia (2020) Omar Khattab and Matei Zaharia. 2020. ColBERT: Efficient and Effective Passage Search via Contextualized Late Interaction over BERT. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval, SIGIR 2020, Virtual Event, China, July 25-30, 2020, Jimmy X. Huang, Yi Chang, Xueqi Cheng, Jaap Kamps, Vanessa Murdock, Ji-Rong Wen, and Yiqun Liu (Eds.). ACM, 39–48. doi:10.1145/3397271.3401075
@@
- Sharma et al. (2024) Rohan Sharma, Kaiyi Ji, Zhiqiang Xu, and Changyou Chen. 2024. AUC-CL: A Batchsize-Robust Framework for Self-Supervised Contrastive Representation Learning. In The Twelfth International Conference on Learning Representations, ICLR 2024, Vienna, Austria, May 7-11, 2024. OpenReview.net. https://openreview.net/forum?id=YgMdDQB09U
@@
- Thakur et al. (2021) Nandan Thakur, Nils Reimers, Andreas Rücklé, Abhishek Srivastava, and Iryna Gurevych. 2021. BEIR: A Heterogeneous Benchmark for Zero-shot Evaluation of Information Retrieval Models. In Proceedings of the Neural Information Processing Systems Track on Datasets and Benchmarks 1, NeurIPS Datasets and Benchmarks 2021, December 2021, virtual, Joaquin Vanschoren and Sai-Kit Yeung (Eds.). https://datasets-benchmarks-proceedings.neurips.cc/paper/2021/hash/65b9eea6e1cc6bb9f0cd2a47751a186f-Abstract-round2.html